OpenStudy (anonymous):

If the spinner is spun 72 times, how many times can you expect it to land on a number smaller than 4. ( there are four numbers that are smaller than 4 and total number of numbers is 10).

OpenStudy (anonymous):

The expected number is the (individual probability) x (# of trials) because the trials are independent. Your total # of trials is 72. Your individual probability is the # of numbers less than 4 over the total number of numbers.

OpenStudy (anonymous):

So 4/10 = x/72

OpenStudy (anonymous):

yes, you have it. Good job.

OpenStudy (anonymous):

So 4/10 = 72/1

OpenStudy (anonymous):

(4/10)(72). Just another way of writing it out.

OpenStudy (anonymous):

Thanks mate!

OpenStudy (anonymous):

You're welcome!